# references.bib

### v1

@@ 3.2 Neural Network Architectures
For this study, we choose ResNet-50, EfficientNet B4 and Swin-B as the representative architectures for the benchmarks. They are detailed in Table 4. ResNet [he2016deep] introduced residual learning by incorporating skip connections, enabling the training of very deep neural networks. Despite being the oldest of the architectures, it is still widely used, due to its training efficiency and accuracy. EfficientNet [tan2019efficientnet] improved image classification on ImageNet by optimizing for network size and computational efficiency using a neural architecture search approach. This was selected as another popular choice of CNN architecture but more explicitly designed for computational efficiency than ResNet. Given, the recent interest in transformers for machine vision, we chose the Swin Transformer [liu2021swin] which improved on Vision Transformer (ViT) by introducing a hierarchical architecture that processes images at multiple scales in an efficient manner. This hierarchical design facilitated capturing both local and global features effectively, addressing the limitations of the single-level processing in ViT.
@@ 3.3 Experiments
  • Training Data Quantity Experiment To assess the significance of the training data quantity for each tested model, we evaluated them on subsets containing 25%, 50%, and 75% of the training data. These subsets were generated through random sampling from all training images across fields; maintaining the original proportion of positive and negative samples from the entire dataset.
@@ 3.4 Training Settings
The models were trained on an Nvidia GeForce RTX 2020 Ti. The initial weights for each model were taken from Pytorch’s version of the models pretrained on ImageNet. Since these models were trained on RGB input, the NIR and red edge channels were set to the pretrained weights for the red channel from these models. Each model was trained for 50 epochs. The model with the highest accuracy on the validation set is selected. The optimizer was stochastic gradient descent (SGD) [robbins1951stochastic]. The initial learning rate was 0.001 and this was decayed to 0.0001 using cosine annealing. The batch size used was 4. Cross entropy loss, also known as log loss, was used as the objective function to train the models. The training settings outlined were determined empirically.
@@ 3.5 Evaluation
In order to give a comprehensive overview of evaluation of a classification model’s performance, we use accuracy, precision, recall and MCC. Since the dataset is mostly balanced, accuracy gives a general overview of the quality of predictions. Accuracy represents that total fraction of correct predictions made by the model and is calculated as follows:
@@
While recall is the fraction of the positive class present in the dataset correctly predicted by the model and is calculated as follows:
@@ 4 Results
ResNet 50 $0.873\pm 0.019$ $0.746\pm 0.019$ $0.863\pm 0.014$ $0.876\pm 0.001$
Efficient B4 $0.830\pm 0.018$ $0.661\pm 0.018$ $0.810\pm 0.014$ $0.847\pm 0.020$
Swin B $0.877\pm 0.025$ $0.756\pm 0.024$ $0.872\pm 0.029$ $0.887\pm 0.035$
@@
Table 5 shows that all models achieved an accuracy greater than 0.8. Swin B has the highest accuracy with 0.877, however, ResNet 50 performs only slightly worse with 0.873. EfficientNet B4 performs less well than the other two models with only 0.83. The models rank the same according to MCC.
@@
Figure 4: Accuracy on each field in the test set. The mean of 5 models from repeated training runs was taken. The error bars represent a range of two standard deviations around the mean.
Figure 5 highlights the importance of NIR in the discrimination of wheat and blackgrass. NIR alone outperforms RGB and RGB + NIR. However, using all spectral bands yields the highest accuracy.
Figure 5: Models trained on different combinations of color channels. Models, identified by the combination of channels used, are ordered by accuracy.
@@
Figure 6: Models with different quantities of training data. The mean of 5 models from repeated training runs was taken. The error bars represent a range of two standard deviations around the mean.
@@ 5 Discussion
Our study shows the need for more research into the challenging visual recognition task of detecting grass weeds in grass crops. The dataset we provide aims to offer an opportunity for further research in this area. Given the scale of the dataset, we hope it offers an opportunity to learn generalised solutions. In order to test a models ability to generalise, the validation and test set are made up of images from unseen fields — that is images from fields not in the training dataset.
Overall, the results in Table 5 show that all models achieve an average accuracy of over 80% using all the training data across all the test fields. We can conclude Swin B is the best performing on average but it is subject to more variability in performance. ResNet 50 is only slightly worse on average while EfficientNet B4 performs worse overall. The precision and recall for all models are similar, with recall only slightly higher for all models, suggesting they are quite balanced, capturing most true positives without too many false positives. By investigating the performance on each field in the test set, we find blackgrass classification accuracy improves as the crops and weeds mature in the late season.
@@ 5.1 Fields
In general, Figure 4, all models performed worse on barley crops compared with wheat crops. This could be the result of fewer images of barley crops being in the dataset. Additionally, all models performed best on the fields where images were taken during the late season. Despite fewer images in the training data of late season crops, these were more accurately classified. This suggests that as the crop and blackgrass mature they become easier to discriminate.
@@ 5.2 Channels
Moreover, the dataset is comprised of multispectral images, and allowed us to test the role different spectral bands could play in discriminating between cereal crops and blackgrass. In line with other research, we find that the addition of NIR reliably improves classification accuracy, shown in Figure 5. For example, Red on its own has an accuracy of 0.772 while Red and NIR has an accuracy of 0.860. Similarly, Red and Blue together have an accuracy of 0.770 but with NIR the accuracy is 0.884. Additionally, we find certain powerful combinations of spectral bands, Red, Blue and NIR being one such example, suggesting that particular discriminating features are present in a subset of spectral bands and that in some cases the addition of other spectral bands adds noise to the classification process. This is also shown by combinations that prove detrimental, like Green, NIR and Red Edge, that performs worse than NIR and Red Edge, and NIR alone. This indicates the features in the other colour channels contribute noise to the detection process. However, it is still the case that using all spectral bands in combination is the most accurate.
@@ 5.3 Training Data Size
Finally, since our dataset is unusually large for an agricultural classification dataset we wanted to determine whether this larger dataset conferred any benefits. We found that up to 6000 images there was an improvement in performance but beyond that there was no performance improvement, as shown in Figure 6. It could be that the variation in the 37 fields in the training set was well represented by the 6000 randomly sampled images and provided no additional discriminating information. While EfficientNet has the lowest accuracy of the approaches tried, it performs the best on a very restricted dataset size.
@@ Acknowledgement
This research was partially supported by a PhD studentship from the University of Lincoln, and by Lincoln Agri-Robotics as part of the Expanding Excellence in England (E3) Programme.
@@ Appendix A Full List of Wheat and Barley Papers by Weed Species
Aethusa cynapium 1 [anderegg2023farm]
Alchemilla arvensis Parsley Piert 2 [anderegg2023farm, xu2023weedsnet]
Alopecurus aequalis Shortawn Foxtail 1 [xu2022multi]
@@
Alopecurus myosuroides Blackgrass 32 [dyrmann2015roboweedsupport, sukumar2016weed, chavan2018agroavnet, binguitcha2019crops, elnemr2019convolutional, jiang2019novel, espejo2020improving, fu2020field, rahman2020performance, skacev2020development, trong2020late, espejo2021combining, espejo2021testing, farkhani2021weed, jensen2021predictive, mounashree2021automated, siddiqui2021neural, trong2021yielding, damilare2022weed, fraccaro2022deep, kumar2022weed, kundur2022deep, mu2022faster, reddy2022weed, srivastava2022role, tara2022machine, bharathi2023plant, dheeraj2023using, guo2023identification, mishra2023comparative, mu2023densenet, rahman2023new]
@@
Amaranthus retroflexus Redroot Pigweed 4 [xu2020recognition, ronay2022effect, xu2022multi, xu2023weedsnet]
@@
Apera spica-venti Loose Silky-bent 30 [dyrmann2015roboweedsupport, chavan2018agroavnet, binguitcha2019crops, elnemr2019convolutional, jiang2019novel, espejo2020improving, fu2020field, rahman2020performance, skacev2020development, trong2020late, espejo2021combining, espejo2021testing, farkhani2021weed, mounashree2021automated, siddiqui2021neural, trong2021yielding, damilare2022weed, kumar2022weed, kundur2022deep, mu2022faster, reddy2022weed, srivastava2022role, tara2022machine, anderegg2023farm, bharathi2023plant, dheeraj2023using, guo2023identification, mishra2023comparative, mu2023densenet, rahman2023new]
Arctotheca calendula Cape Dandelion 1 [le2021detecting]
Avena fatua Wild Oat 1 [shahbazi2021assessing]
Bromus japonicus Japanese brome 2 [xu2022multi, xu2023weedsnet]
@@
Capsella bursa-pastoris Shepherd’s Purse 36 [dyrmann2015roboweedsupport, dyrmann2017roboweedsupport, chavan2018agroavnet, dyrmann2018using, binguitcha2019crops, elnemr2019convolutional, jiang2019novel, espejo2020improving, fu2020field, rahman2020performance, skacev2020development, trong2020late, xu2020recognition, espejo2021combining, espejo2021testing, farkhani2021weed, mounashree2021automated, siddiqui2021neural, trong2021yielding, damilare2022weed, kumar2022weed, kundur2022deep, mu2022faster, reddy2022weed, srivastava2022role, tara2022machine, xu2022multi, zhuang2022evaluation, anderegg2023farm, bharathi2023plant, dheeraj2023using, guo2023identification, liu2023semi, mishra2023comparative, mu2023densenet, rahman2023new]
Centaurea cyanus Cornflower 1 [anderegg2023farm]
@@
Chenopodium album Fat Hen 35 [dyrmann2015roboweedsupport, dyrmann2017roboweedsupport, chavan2018agroavnet, dyrmann2018using, wang2018low, binguitcha2019crops, elnemr2019convolutional, jiang2019novel, espejo2020improving, fu2020field, rahman2020performance, skacev2020development, trong2020late, wang2020transfer, espejo2021combining, espejo2021testing, farkhani2021weed, mounashree2021automated, siddiqui2021neural, trong2021yielding, damilare2022weed, kumar2022weed, kundur2022deep, mu2022faster, reddy2022weed, srivastava2022role, tara2022machine, alirezazadeh2023comparative, anderegg2023farm, bharathi2023plant, dheeraj2023using, guo2023identification, mishra2023comparative, mu2023densenet, rahman2023new]
Cerastium fontanum Common Mouse Ear Chickweed 1 [anderegg2023farm]
Chloris cucullata Hooded Windmill Grass 1 [saqib2023towards]
@@
Cirsium arvense Creeping Thistle 5 [franco2018automatic, rasmussen2019pre, rasmussen2021pre, haq2023weed, saqib2023towards, thomas2023weakly]
@@
Consolida regalis Forking Larkspur 3 [almalky2022efficient, almalky2023deep, almalky2023real]
@@
Convolvulus arvensis Field Bindweed 3 [raptis2022multimodal, anderegg2023farm, saqib2023towards]
Dactylis glomerata Cat Grass 1 [saqib2023towards]
Echinochloa crus-galli Cockspur Grass 2 [xu2022multi, xu2023weedsnet]
Elymus repens Common Couch 1 [thomas2023weakly]
Equisetum arvense Horsetail 1 [zou2021modified]
Eschscholzia californica California poppy 1 [saqib2023towards]
Fallopia convolvulus Black Bindweed 1 [anderegg2023farm]
Fumaria officinalis Common fumitory 1 [anderegg2023farm]
@@
Galium aparine Cleavers 34 [dyrmann2015roboweedsupport, dyrmann2017roboweedsupport, chavan2018agroavnet, dyrmann2018using, binguitcha2019crops, elnemr2019convolutional, jiang2019novel, espejo2020improving, fu2020field, rahman2020performance, skacev2020development, trong2020late, espejo2021combining, espejo2021testing, farkhani2021weed, mounashree2021automated, siddiqui2021neural, trong2021yielding, damilare2022weed, kumar2022weed, kundur2022deep, mu2022faster, reddy2022weed, srivastava2022role, tara2022machine, zhuang2022evaluation, anderegg2023farm, bharathi2023plant, dheeraj2023using, guo2023identification, liu2023semi, mishra2023comparative, mu2023densenet, rahman2023new]
Galium odoratum Sweet Woodruff 1 [liu2023semi]
@@
Geranium pusillum Small-flowered Cranesbill 30 [dyrmann2017roboweedsupport, chavan2018agroavnet, dyrmann2018using, binguitcha2019crops, elnemr2019convolutional, jiang2019novel, espejo2020improving, fu2020field, rahman2020performance, skacev2020development, trong2020late, espejo2021combining, espejo2021testing, farkhani2021weed, mounashree2021automated, siddiqui2021neural, trong2021yielding, damilare2022weed, kumar2022weed, kundur2022deep, mu2022faster, reddy2022weed, srivastava2022role, tara2022machine, bharathi2023plant, dheeraj2023using, guo2023identification, mishra2023comparative, mu2023densenet, rahman2023new]
Humulus japonicus Japanese Hops 2 [wang2018low, wang2020transfer]
Lamium purpureum Red Deadnettle 2 [alirezazadeh2023comparative, anderegg2023farm]
@@
Lolium multiflorum Italian Ryegrass 2 [jensen2021predictive, anderegg2023farm]
@@
Lolium perenne Perennial Ryegrass 3 [dyrmann2017roboweedsupport, dyrmann2018using, saqib2023towards]
Lipandra polysperma Manyseed Goosefoot 1 [anderegg2023farm]
Lithospermum arvense Corn Gromwell 1 [alirezazadeh2023comparative]
Matricaria chamomilla Chamomile 1 [anderegg2023farm]
Matricaria discoidea Pineappleweed 2 [sukumar2016weed, pflanz2018weed]
@@
Myosotis arvensis Field Forget-Me-Not 3 [dyrmann2017roboweedsupport, dyrmann2018using, anderegg2023farm]
@@
Papaver rhoeas Common Poppy 3 [pflanz2018weed, anderegg2023farm, perez2023early]
@@
Persicaria maculosa Redshank 3 [dyrmann2017roboweedsupport, dyrmann2018using, anderegg2023farm]
Phalaris paradoxa Awned Canarygrass 1 [jabir2021accuracy]
Poa annua Annual Bluegrass 3 [xu2020recognition, xu2022multi, anderegg2023farm]
Poa pratensis Kentucky bluegrass 1 [xu2023weedsnet]
Polygonum aviculare Common Knotweed 1 [anderegg2023farm]
Portulaca oleracea Purslane 1 [raptis2022multimodal]
Raphanus raphanistrum Wild Radish 2 [le2021detecting, shahbazi2021assessing]
Rumex crispus Curly Dock 2 [el2022metaheuristic, pp2023metaheuristic]
Rumex obtusifolius Broadleaf Dock 2 [el2022metaheuristic, pp2023metaheuristic]
Setaria adhaerens 1 [ronay2022effect]
Setaria viridis Green Foxtail 1 [zou2021modified]
@@
Sinapis arvensis Charlock 31 [dyrmann2015roboweedsupport, dyrmann2017roboweedsupport, chavan2018agroavnet, dyrmann2018using, binguitcha2019crops, elnemr2019convolutional, jiang2019novel, espejo2020improving, fu2020field, rahman2020performance, skacev2020development, trong2020late, espejo2021combining, espejo2021testing, farkhani2021weed, mounashree2021automated, siddiqui2021neural, trong2021yielding, damilare2022weed, kumar2022weed, kundur2022deep, mu2022faster, reddy2022weed, srivastava2022role, tara2022machine, bharathi2023plant, dheeraj2023using, guo2023identification, mishra2023comparative, mu2023densenet, rahman2023new]
Sorghum halepense Johnson grass 1 [raptis2022multimodal]
@@
Solanum nigrum Black Nightshade 4 [dyrmann2017roboweedsupport, dyrmann2018using, ronay2022effect, anderegg2023farm]
Spergula arvensis Corn Spurry 1 [alirezazadeh2023comparative]

### v2

@@ 3.2 Neural Network Architectures
For this study, we choose ResNet-50, EfficientNet B4 and Swin-B as the representative architectures for the benchmarks. They are detailed in Table 4. ResNet he2016deep introduced residual learning by incorporating skip connections, enabling the training of very deep neural networks. Despite being the oldest of the architectures, it is still widely used, due to its training efficiency and accuracy. EfficientNet tan2019efficientnet improved image classification on ImageNet by optimizing for network size and computational efficiency using a neural architecture search approach. This was selected as another popular choice of CNN architecture but more explicitly designed for computational efficiency than ResNet. Given, the recent interest in transformers for machine vision, we chose the Swin Transformer liu2021swin which improved on Vision Transformer (ViT) by introducing a hierarchical architecture that processes images at multiple scales in an efficient manner. This hierarchical design facilitated capturing both local and global features effectively, addressing the limitations of the single-level processing in ViT.
@@ 3.3 Experiments
  • Training Data Quantity Experiment To assess the significance of the training data quantity for each tested model, we evaluated them on subsets containing 1%, 3%, 6%, 12%, 25%, 50%, and 75% of the training data. These subsets were generated through random sampling from all training images across fields; maintaining the original proportion of positive and negative samples from the entire dataset.
@@ 3.4 Training Settings
The models were trained on an Nvidia GeForce RTX 2020 Ti. The initial weights for each model were taken from Pytorch’s version of the models pretrained on ImageNet. Since these models were trained on RGB input, the NIR and red edge channels were set to the pretrained weights for the red channel from these models. Each model was trained for 50 epochs. The model with the highest accuracy on the validation set is selected. The optimizer was stochastic gradient descent (SGD) robbins1951stochastic . The initial learning rate was 0.001 and this was decayed to 0.0001 using cosine annealing. The batch size used was 8. Cross entropy loss, also known as log loss, was used as the objective function to train the models. The training settings outlined were determined empirically.
@@ 3.5 Evaluation
In order to give a comprehensive overview of evaluation of a classification model’s performance, we use four standard classification metrics: accuracy, precision, recall and MCC. Since the dataset is mostly balanced, accuracy gives a general overview of the quality of predictions. Accuracy represents that total fraction of correct predictions made by the model and is calculated as follows:
@@
Recall is the fraction of the positive class present in the dataset correctly predicted by the model and is calculated as follows:
@@ 4 Results
ResNet 50 $0.866\pm 0.026$ $0.855\pm 0.010$ $0.851\pm 0.044$ $0.733\pm 0.050$
Efficient B4 $0.833\pm 0.010$ $0.869\pm 0.036$ $0.799\pm 0.050$ $0.670\pm 0.019$
Swin B $0.896\pm 0.024$ $0.912\pm 0.051$ $0.886\pm 0.028$ $0.794\pm 0.048$
@@
Table 5 shows that all models achieved an accuracy greater than 0.8. Swin B has the highest accuracy with 0.896, however, ResNet 50 performs only slightly worse with 0.866. EfficientNet B4 performs less well than the other two models with only 0.833.
@@
Figure 4: The accuracy of each model—ResNet50 (blue), EfficientNet B4 (orange), and Swin B (green)—on each field in the test set. The mean of 5 models from repeated training runs was taken. The error bars represent a range of two standard deviations around the mean.
Figure 5 highlights the importance of NIR in the discrimination of wheat and blackgrass.
Figure 5: Models trained on different combinations of color channels. Models, identified by the combination of channels used, are ordered by accuracy. The mean of 3 models from repeated training runs was taken. The error bars represent a range of two standard deviations around the mean.
@@
Figure 6: The accuracy of each model—ResNet50 (blue), EfficientNet B4 (orange), and Swin B (green)—with different quantities of training data. The mean of 5 models from repeated training runs was taken. The coloured bands represent a range of two standard deviations around the mean.
@@ 5 Discussion
Our analysis of the existing literature showed the need for more research into the challenging visual recognition task of recognising grass weeds in grass crops. The dataset we provide aims to offer an opportunity for further research in this area. Given the scale of the dataset, we hope it offers an opportunity to learn generalised solutions. In order to test a model’s ability to generalise, the validation and test set are made up of images from unseen fields — that is images from fields not in the training dataset.
Overall, the results in Table 5 show that all models achieve an average accuracy of over 80% using all the training data across all the test fields. We can conclude Swin B is the best performing on average. ResNet 50 is only slightly worse on average while EfficientNet B4 performs worse overall. All models exhibit high precision, indicating few false positives, but lower recall suggests more false negatives–that is, more blackgrass instances are being missed.
@@ 5.1 Fields
In general, Figure 4 shows all models performed worse on barley crops compared with wheat crops. This could be the result of fewer images of barley crops being in the dataset. Additionally, all models performed best on the fields where images were taken during the late season. Despite fewer images in the training data of late season crops, these were more accurately classified. This suggests that as the crop and blackgrass mature they become easier to discriminate.
@@ 5.2 Channels
The dataset is comprised of multispectral images, and allowed us to test the role different spectral bands could play in discriminating between cereal crops and blackgrass. In line with other research, we find that the addition of NIR reliably improves classification accuracy sa2018weednet ; wang2020semantic , shown in Figure 5. For example, Red on its own has an accuracy of 0.741 while Red and NIR has an accuracy of 0.859. Similarly, Red and Blue together have an accuracy of 0.776 but with NIR the accuracy is 0.873. Furthermore, the 14 most accurate models all have NIR as an input.
Some powerful subsets like Red, Red Edge and NIR outperform the use of all channels on average, suggesting that particular discriminating features are present in a subset of spectral bands and that in some cases the addition of other spectral bands adds noise to the classification process.
@@ 5.3 Training Data Size
Finally, since our dataset is unusually large for an agricultural classification dataset we wanted to determine whether this larger dataset conferred any benefits. We found that up to 6000 images there was an improvement in performance but beyond that the performance improvement was small, as shown in 6. It could be that the variation in the 37 fields in the training set was well represented by the 6000 randomly sampled images and additional images provided minimal additional discriminating information. While EfficientNet has the lowest accuracy of the approaches tried, it performs the best on a very restricted dataset size.
5.4 Limitations
While many popular crop and weed datasets have bounding box labels or pixel-level labels, our dataset only has image-level labels. This is due to the fact that to the classification of blackgrass could often not be made by looking at the image, especially during the early season, even by an expert observer. Instead, a plant’s anatomy would need to be carefully examined in-person to make such a classification. However, we acknowledge that a weed management system trained on these images may have limited precision, as it can only recognise blackgrass at the image level.
At present, this work focuses on recognising the presence of black- grass, however, later work will investigate blackgrass density classification.
@@ 6 Conclusion
Data Availability Statement
The dataset used in this study will be made available for non-commercial use under the Creative Commons Attribution-NonCommercial-ShareAlike (CC BY-NC-SA) license. Access to the dataset will be granted upon request, subject to agreement with the terms of the license.
@@ Acknowledgement
This research was partially supported by a PhD studentship from the University of Lincoln, and by Lincoln Agri-Robotics as part of the Expanding Excellence in England (E3) Programme. We would like to acknowledge support from the UKRI Research England Ceres Agri-tech grant for funding the collection of the data set and initial testing in project Ceres:A13722 IC8P1; Autonomous Black-Grass Detection. We would also like to thanks the land managers that helped us identify fields and black-grass patches, and gave us access for image collection.
@@ Appendix A Full List of Wheat and Barley Papers by Weed Species
Aethusa cynapium 1 anderegg2023farm
Alchemilla arvensis Parsley Piert 2 anderegg2023farm ; xu2023weedsnet
Alopecurus aequalis Shortawn Foxtail 1 xu2022multi
@@
Alopecurus myosuroides Blackgrass 32 dyrmann2015roboweedsupport ; sukumar2016weed ; chavan2018agroavnet ; binguitcha2019crops ; elnemr2019convolutional ; jiang2019novel ; espejo2020improving ; fu2020field ; rahman2020performance ; skacev2020development ; trong2020late ; espejo2021combining ; espejo2021testing ; farkhani2021weed ; jensen2021predictive ; mounashree2021automated ; siddiqui2021neural ; trong2021yielding ; damilare2022weed ; fraccaro2022deep ; kumar2022weed ; kundur2022deep ; mu2022faster ; reddy2022weed ; srivastava2022role ; tara2022machine ; bharathi2023plant ; dheeraj2023using ; guo2023identification ; mishra2023comparative ; mu2023densenet ; rahman2023new
@@
Amaranthus retroflexus Redroot Pigweed 4 xu2020recognition ; ronay2022effect ; xu2022multi ; xu2023weedsnet
@@
Apera spica-venti Loose Silky-bent 30 dyrmann2015roboweedsupport ; chavan2018agroavnet ; binguitcha2019crops ; elnemr2019convolutional ; jiang2019novel ; espejo2020improving ; fu2020field ; rahman2020performance ; skacev2020development ; trong2020late ; espejo2021combining ; espejo2021testing ; farkhani2021weed ; mounashree2021automated ; siddiqui2021neural ; trong2021yielding ; damilare2022weed ; kumar2022weed ; kundur2022deep ; mu2022faster ; reddy2022weed ; srivastava2022role ; tara2022machine ; anderegg2023farm ; bharathi2023plant ; dheeraj2023using ; guo2023identification ; mishra2023comparative ; mu2023densenet ; rahman2023new
Arctotheca calendula Cape Dandelion 1 le2021detecting
Avena fatua Wild Oat 1 shahbazi2021assessing
Bromus japonicus Japanese brome 2 xu2022multi ; xu2023weedsnet
@@
Capsella bursa-pastoris Shepherd’s Purse 36 dyrmann2015roboweedsupport ; dyrmann2017roboweedsupport ; chavan2018agroavnet ; dyrmann2018using ; binguitcha2019crops ; elnemr2019convolutional ; jiang2019novel ; espejo2020improving ; fu2020field ; rahman2020performance ; skacev2020development ; trong2020late ; xu2020recognition ; espejo2021combining ; espejo2021testing ; farkhani2021weed ; mounashree2021automated ; siddiqui2021neural ; trong2021yielding ; damilare2022weed ; kumar2022weed ; kundur2022deep ; mu2022faster ; reddy2022weed ; srivastava2022role ; tara2022machine ; xu2022multi ; zhuang2022evaluation ; anderegg2023farm ; bharathi2023plant ; dheeraj2023using ; guo2023identification ; liu2023semi ; mishra2023comparative ; mu2023densenet ; rahman2023new
Centaurea cyanus Cornflower 1 anderegg2023farm
@@
Chenopodium album Fat Hen 35 dyrmann2015roboweedsupport ; dyrmann2017roboweedsupport ; chavan2018agroavnet ; dyrmann2018using ; wang2018low ; binguitcha2019crops ; elnemr2019convolutional ; jiang2019novel ; espejo2020improving ; fu2020field ; rahman2020performance ; skacev2020development ; trong2020late ; wang2020transfer ; espejo2021combining ; espejo2021testing ; farkhani2021weed ; mounashree2021automated ; siddiqui2021neural ; trong2021yielding ; damilare2022weed ; kumar2022weed ; kundur2022deep ; mu2022faster ; reddy2022weed ; srivastava2022role ; tara2022machine ; alirezazadeh2023comparative ; anderegg2023farm ; bharathi2023plant ; dheeraj2023using ; guo2023identification ; mishra2023comparative ; mu2023densenet ; rahman2023new
Cerastium fontanum Common Mouse Ear Chickweed 1 anderegg2023farm
Chloris cucullata Hooded Windmill Grass 1 saqib2023towards
@@
Cirsium arvense Creeping Thistle 5 franco2018automatic ; rasmussen2019pre ; rasmussen2021pre ; haq2023weed ; saqib2023towards ; thomas2023weakly
@@
Consolida regalis Forking Larkspur 3 almalky2022efficient ; almalky2023deep ; almalky2023real
@@
Convolvulus arvensis Field Bindweed 3 raptis2022multimodal ; anderegg2023farm ; saqib2023towards
Dactylis glomerata Cat Grass 1 saqib2023towards
Echinochloa crus-galli Cockspur Grass 2 xu2022multi ; xu2023weedsnet
Elymus repens Common Couch 1 thomas2023weakly
Equisetum arvense Horsetail 1 zou2021modified
Eschscholzia californica California poppy 1 saqib2023towards
Fallopia convolvulus Black Bindweed 1 anderegg2023farm
Fumaria officinalis Common fumitory 1 anderegg2023farm
@@
Galium aparine Cleavers 34 dyrmann2015roboweedsupport ; dyrmann2017roboweedsupport ; chavan2018agroavnet ; dyrmann2018using ; binguitcha2019crops ; elnemr2019convolutional ; jiang2019novel ; espejo2020improving ; fu2020field ; rahman2020performance ; skacev2020development ; trong2020late ; espejo2021combining ; espejo2021testing ; farkhani2021weed ; mounashree2021automated ; siddiqui2021neural ; trong2021yielding ; damilare2022weed ; kumar2022weed ; kundur2022deep ; mu2022faster ; reddy2022weed ; srivastava2022role ; tara2022machine ; zhuang2022evaluation ; anderegg2023farm ; bharathi2023plant ; dheeraj2023using ; guo2023identification ; liu2023semi ; mishra2023comparative ; mu2023densenet ; rahman2023new
Galium odoratum Sweet Woodruff 1 liu2023semi
@@
Geranium pusillum Small-flowered Cranesbill 30 dyrmann2017roboweedsupport ; chavan2018agroavnet ; dyrmann2018using ; binguitcha2019crops ; elnemr2019convolutional ; jiang2019novel ; espejo2020improving ; fu2020field ; rahman2020performance ; skacev2020development ; trong2020late ; espejo2021combining ; espejo2021testing ; farkhani2021weed ; mounashree2021automated ; siddiqui2021neural ; trong2021yielding ; damilare2022weed ; kumar2022weed ; kundur2022deep ; mu2022faster ; reddy2022weed ; srivastava2022role ; tara2022machine ; bharathi2023plant ; dheeraj2023using ; guo2023identification ; mishra2023comparative ; mu2023densenet ; rahman2023new
Humulus japonicus Japanese Hops 2 wang2018low ; wang2020transfer
Lamium purpureum Red Deadnettle 2 alirezazadeh2023comparative ; anderegg2023farm
@@
Lolium multiflorum Italian Ryegrass 2 jensen2021predictive ; anderegg2023farm
@@
Lolium perenne Perennial Ryegrass 3 dyrmann2017roboweedsupport ; dyrmann2018using ; saqib2023towards
Lipandra polysperma Manyseed Goosefoot 1 anderegg2023farm
Lithospermum arvense Corn Gromwell 1 alirezazadeh2023comparative
Matricaria chamomilla Chamomile 1 anderegg2023farm
Matricaria discoidea Pineappleweed 2 sukumar2016weed ; pflanz2018weed
@@
Myosotis arvensis Field Forget-Me-Not 3 dyrmann2017roboweedsupport ; dyrmann2018using ; anderegg2023farm
@@
Papaver rhoeas Common Poppy 3 pflanz2018weed ; anderegg2023farm ; perez2023early
@@
Persicaria maculosa Redshank 3 dyrmann2017roboweedsupport ; dyrmann2018using ; anderegg2023farm
Phalaris paradoxa Awned Canarygrass 1 jabir2021accuracy
Poa annua Annual Bluegrass 3 xu2020recognition ; xu2022multi ; anderegg2023farm
Poa pratensis Kentucky bluegrass 1 xu2023weedsnet
Polygonum aviculare Common Knotweed 1 anderegg2023farm
Portulaca oleracea Purslane 1 raptis2022multimodal
Raphanus raphanistrum Wild Radish 2 le2021detecting ; shahbazi2021assessing
Rumex crispus Curly Dock 2 el2022metaheuristic ; pp2023metaheuristic
Rumex obtusifolius Broadleaf Dock 2 el2022metaheuristic ; pp2023metaheuristic
Setaria adhaerens 1 ronay2022effect
Setaria viridis Green Foxtail 1 zou2021modified
@@
Sinapis arvensis Charlock 31 dyrmann2015roboweedsupport ; dyrmann2017roboweedsupport ; chavan2018agroavnet ; dyrmann2018using ; binguitcha2019crops ; elnemr2019convolutional ; jiang2019novel ; espejo2020improving ; fu2020field ; rahman2020performance ; skacev2020development ; trong2020late ; espejo2021combining ; espejo2021testing ; farkhani2021weed ; mounashree2021automated ; siddiqui2021neural ; trong2021yielding ; damilare2022weed ; kumar2022weed ; kundur2022deep ; mu2022faster ; reddy2022weed ; srivastava2022role ; tara2022machine ; bharathi2023plant ; dheeraj2023using ; guo2023identification ; mishra2023comparative ; mu2023densenet ; rahman2023new
Sorghum halepense Johnson grass 1 raptis2022multimodal
@@
Solanum nigrum Black Nightshade 4 dyrmann2017roboweedsupport ; dyrmann2018using ; ronay2022effect ; anderegg2023farm
Spergula arvensis Corn Spurry 1 alirezazadeh2023comparative